Work Project Report

Jaroslav Vazny Masaryk University, Czech Republic

            jaroslav.vazny@gmail.com

August 24, 2010

Abstract

This is my project report for summer student program in 2010. I was working on improving beam emittance and luminosity of the BSD (Beam Delivery System) part of future CLIC (Compact Linear Collider).

1 Project Targets

CLIC specification (3 TeV) Initial

2 Results

Parameter Designed(3TeV) Initial Final Improvements height

3 Implementation

3.1 Used Tools

madx placet shell-scripts python

3.2 Version 1 3.3 Version 2

4 Conclusion and Recommandations

4.1 Automatization 4.2 Version Control